# USTC-ICTS/PCFT-25-53

### v1

@@ 1 Introduction
The cosmological constant comes from vacuum energy in quantum field theory. In generic scenarios, this calculation is divergent so naturally it should be of the order of the UV cut-off Planck energy. However, the observed value is about $10^{-122}$ of the naive expectation. This is the famous cosmological constant problem. We should note that before the discovery of accelerated expansion by supernovae experiments in the late 1990s, it was commonly believed that the cosmological constant is probably zero and many mechanisms were proposed to explain this exact cancellation. A small but non-zero value seems more difficult to explain. However this also provides an opportunity that if we can find a natural way to estimate the right order of magnitude of the cosmological constant, the arguments would be more quantitative and convincing.
There have been many proposed solutions of the cosmological constant problem, but there is no universally accepted consensus at the moment. We will consider this problem from the perspective of entropy. For an observer in de Sitter space, there is a cosmological event horizon where a light signal sent by the observer can not reach even after an infinite travel time, due to the accelerated expansion of the space. In 1970’s, Gibbons and Hawking proposed that similar to the Bekenstein-Hawking black hole entropy [1, 2], the cosmological horizon also measures the entropy of degrees of freedom beyond the horizon [3]. The Gibbons-Hawking entropy is proportional to the area of the cosmological horizon, and is inversely proportional to the cosmological constant. So for our universe, this entropy is about $10^{122}$. (In this paper we use the convention $\hbar=c=k_{B}=1$ and entropy is a dimensionless quantity.) If one can find a natural way to estimate the entropy of our universe, we can then also obtain an estimate for the cosmological constant.
@@ 2.1 Gibbons-Hawking entropy
We note that the calculation of Gibbons-Hawking entropy is not affected by possible extra compact dimensions and their sizes. Suppose we compactify from $d$ dimensions to 4 dimensions, then the length scales of gravity are related $l_{d}^{d-2}=V_{d-4}l_{p}^{2}$, where $V_{d-4}$ is the compactification volume. The entropy is the same for the $d$-dimensional calculation $S=\frac{V_{d-4}A}{4l_{d}^{d-2}}=\frac{A}{4l_{p}^{2}}$, where $A$ is the 4d horizon area.
@@ 2.2 Other entropies
There have been some estimates of various components of the observable universe in astrophysics, see e.g. [12, 13]. In generic scenarios, the dominant contribution to entropy comes from supermassive black holes (SMBH), which reside at the center of most galaxies, including our own Milky Way. A typical SMBH may have a mass of $10^{8}$ times that of the Sun and an entropy of $10^{93}$. It is estimated that there are no more than $10^{12}$ galaxies in the observable universe. So the total entropy of SMBHs is about $10^{105}$. How SMBHs formed so early and grew so huge is still one of the big open questions in astrophysics.
@@ 3 A universal finite upper bound for entropy?
We now describe a physical process to generate entropy. We start with a pure state, which has zero von Neumann entropy. The initial state can then undergo a quantum evolution which we assume to be unitary. A unitary transformation does not change the von Neumann entropy. In order to have a finite entropy, we allow for an external observer to make a measurement. In the usual interpretation of quantum measurements, the quantum state is projected into an orthogonal basis of the Hilbert space, and we have a probability distribution for the outcomes of the measurement. Suppose the external observer does not reveal the result of the measurement, then we have a mixed state with finite entropy. This is the usual decoherence process where the environment plays the role of an external observer.
@@
where $p_{n}$’s form a probability distribution. A simple example is the symmetric random walk, where one moves left or right by one unit with equal probability, i.e. $p_{1}=p_{-1}=\frac{1}{2}$ and all other $p_{n}$’s vanish. Suppose we start with the state $|0\rangle$ and perform the unitarity transformation and an external measurement in the basis $|n\rangle,n\in\mathbb{Z}$, then we have a mixed state $\sum_{n=-\infty}^{\infty}p_{n}|n\rangle\langle n|$, which can be equivalently represented as a random variable taking values in integers with the prescribed probability. If we repeat this procedure $N$ times, we get a mix state represented by the sum of $N$ such random variables. When $N$ is large, according to the Central Limit Theorem, the probability is well approximated by a normal distribution whose variance is proportional to $N$. We can compute the entropy with the continuous normal distribution $f(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{x^{2}}{2\sigma^{2}}}$ as a good approximation if $\sigma$ is large
@@
Perhaps an easier problem is to first determine whether the $\lim_{g\rightarrow\infty}p_{n.m}(g)$ is zero or non-zero. Here we describe a proposition which may be useful if this limit is zero. For simplicity we omit the initial state index $m$. Proposition: suppose $\lim_{g\rightarrow\infty}p_{n}(g)=0$ uniformly in $n$, then the entropy $S(g)=-\sum_{n}p_{n}(g)\log p_{n}(g)$ tends to infinity as $g\rightarrow\infty$. This proposition applies for example to the thermal entropies in [7]. The proof is not difficult. For any large natural number $N$, we can find a $g_{0}$ such that when $g>g_{0}$, we have $p_{n}(g)<\frac{1}{N}$ for any $n$. We note that the function $f(x)=-x\log(x)$ is a concave function for $x\geq 0$ because its second derivative is always negative. So the entropy is decreased if we move $N$ of the $p_{n}$’s to $\frac{1}{N}$ and the remaining ones to zero, keeping the sum fixed $\sum_{n}p_{n}=1$. In this case the entropy is $\log(N)$, which tends to infinity as $N\rightarrow\infty$.
@@
The proposition can be slightly generalized. Suppose $\lim_{g\rightarrow\infty}p_{n}(g)=\tilde{p}_{n}$ uniformly in $n$, only a finite number of $\tilde{p}_{n}$’s are non-zero, and $\sum_{n}\tilde{p}_{n}<1$, then the entropy also tends to infinity as $g\rightarrow\infty$. The proof is similar.

### v2

@@ 1 Introduction
The cosmological constant $\Lambda$ comes from vacuum energy in quantum field theory. In generic scenarios, this calculation is divergent so naturally it should be of the order of the UV cut-off Planck energy. However, the observed value is about $10^{-122}$ of the naive expectation. This is the famous cosmological constant problem. We should note that before the discovery of accelerated expansion by supernovae experiments in the late 1990s, it was commonly believed that $\Lambda$ is probably zero and many mechanisms were proposed to explain this exact cancellation. A small but non-zero value seems more difficult to explain. However this also provides an opportunity that if we can find a natural way to estimate the right order of magnitude of the cosmological constant, the arguments would be more quantitative and convincing.
There have been many proposed solutions of the cosmological constant problem, but there is no universally accepted consensus at the moment. We will consider this problem from the perspective of entropy. In this paper we put aside the possibility of $\Lambda$ being zero or negative, and focus on the case of asymptotic de Sitter space with a positive $\Lambda$. In this case, there is a cosmological event horizon where a light signal sent by an observer cannot reach even after an infinite travel time, due to the accelerated expansion of the space. In 1970s, Gibbons and Hawking proposed that similar to the Bekenstein-Hawking black hole entropy [1, 2], the cosmological horizon also measures the entropy of degrees of freedom beyond the horizon [3]. The Gibbons-Hawking entropy is proportional to the area of the cosmological horizon, and is inversely proportional to the cosmological constant $\Lambda$. So for our universe, this entropy is about $10^{122}$. (In this paper we use the convention $\hbar=c=k_{B}=1$ and entropy is a dimensionless quantity.) If one can find a natural way to estimate the entropy of our universe, we can then also obtain an estimate for the cosmological constant.
@@ 2.1 Gibbons-Hawking entropy
Before late 1990s, some occasional pioneering works did seriously consider a non-zero $\Lambda$ for our universe. Notably, Weinberg’s anthropic argument [10] essentially predicted the cosmological coincidence, i.e. the current energy densities of $\Lambda$ and matter are comparable $\Omega_{\Lambda}\sim\Omega_{m}$. If $\Omega_{\Lambda}$ is order one, then we see that the order of magnitude of $\Lambda$ is determined by the observed Hubble constant. It is a much more difficult problem to provide a purely theoretical derivation of this absolute magnitude $10^{-122}$ in Planck unit.
We provide some more general discussions about Gibbons-Hawking entropy. We note that the calculation is not affected by possible extra compact dimensions and their sizes. Suppose we compactify from $d$ dimensions to 4 dimensions, then the length scales of gravity are related $l_{d}^{d-2}=V_{d-4}l_{p}^{2}$, where $V_{d-4}$ is the compactification volume. The entropy is the same for the $d$-dimensional calculation $S=\frac{V_{d-4}A}{4l_{d}^{d-2}}=\frac{A}{4l_{p}^{2}}$, where $A$ is the 4d horizon area.
@@ 2.2 Other entropies
There have been some estimates for the entropies of various components of the observable universe in astrophysics, see e.g. [13, 14]. In generic scenarios, the dominant contribution to entropy comes from supermassive black holes (SMBH), which reside at the center of most galaxies, including our own Milky Way. A typical SMBH may have a mass of $10^{8}$ times that of the Sun and an entropy of $10^{93}$. It is estimated that there are no more than $10^{12}$ galaxies in the observable universe. So the total entropy of SMBHs is about $10^{105}$. How SMBHs formed so early and grew so huge is still one of the big open questions in astrophysics.
@@ 3 A universal finite upper bound for entropy?
We now describe a physical process to generate entropy. We start with a pure state, which has zero von Neumann entropy. The initial state can then undergo a quantum evolution which we assume to be unitary. A unitary transformation does not change the von Neumann entropy. In order to have a finite entropy, we allow for an external observer to make a measurement. In the usual interpretation of quantum measurements, the quantum state is projected into an orthogonal basis of the Hilbert space, and we have a probability distribution for the outcomes of the measurement. Suppose the external observer does not reveal the result of the measurement, then we have a mixed state with finite entropy. This is the usual decoherence process where the environment plays the role of an external observer. This can be also understood as an entanglement entropy between the quantum state and the observer.
@@
where $p_{n}$’s form a probability distribution. A simple example is the symmetric random walk, where one moves left or right by one unit with equal probability, i.e. $p_{1}=p_{-1}=\frac{1}{2}$ and all other $p_{n}$’s vanish. Suppose we start with the state $|0\rangle$ and perform the unitarity transformation and an external measurement in the basis $|n\rangle,n\in\mathbb{Z}$, then we have a mixed state $\sum_{n=-\infty}^{\infty}p_{n}|n\rangle\langle n|$, which can be equivalently represented as a random variable taking values in integers with the prescribed probability. If we repeat this procedure $N$ times, we get a mixed state represented by the sum of $N$ such random variables. When $N$ is large, according to the Central Limit Theorem, the probability is well approximated by a normal distribution whose variance is proportional to $N$. We can compute the entropy with the continuous normal distribution $f(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{x^{2}}{2\sigma^{2}}}$ as a good approximation if $\sigma$ is large
@@
Perhaps an easier problem is to first understand the asymptotic behavior of $p_{n.m}(g)$ as $g\rightarrow\infty$. Here we prove a proposition which provides a relevant general criterion for the (un)boundedness of entropy if the limit exists. For simplicity we omit the initial state index $m$ and first prove an easier lemma. Lemma: Suppose $\lim_{g\rightarrow\infty}p_{n}(g)=0$ uniformly in $n$, then the entropy $S(g)=-\sum_{n}p_{n}(g)\log p_{n}(g)$ tends to infinity as $g\rightarrow\infty$. This lemma applies for example to the thermal entropies in [7]. The proof is not difficult. For any large natural number $N$, we can find a $g_{0}$ such that when $g>g_{0}$, we have $p_{n}(g)<\frac{1}{N}$ for any $n$. We note that the function $f(x)=-x\log(x)$ is a concave function for $x\geq 0$ because its second derivative is always negative. So the entropy is decreased if we move the $N$ largest $p_{n}(g)$’s to $\frac{1}{N}$ and the remaining ones to zero, keeping the sum fixed $\sum_{n}p_{n}(g)=1$. In this case the entropy is $\log(N)$, which tends to infinity as $N\rightarrow\infty$.
@@
We now prove the stronger proposition. Suppose $\lim_{g\rightarrow\infty}p_{n}(g)=\tilde{p}_{n}$. It is easy to see $0\leq\sum_{n}\tilde{p}_{n}\leq 1$. The sum can be smaller than one because the infinite sum and limit does not necessarily commute. Proposition: Suppose $\lim_{g\rightarrow\infty}p_{n}(g)=\tilde{p}_{n}$ uniformly in $n$ and $\sum_{n}\tilde{p}_{n}<1$, then the entropy also tends to infinity as $g\rightarrow\infty$. Proof: We choose a fixed number $\delta$ in the range $0<\delta<1-\sum_{n}\tilde{p}_{n}$. For any large natural number $N$, since the sum $\sum_{n}\tilde{p}_{n}$ is convergent, only a finite number of $\tilde{p}_{n}$’s are larger than $\frac{\delta}{2N}$. We find a number $n_{0}$ such that when $|n|>n_{0}$, we have $\tilde{p}_{n}<\frac{\delta}{2N}$. Since the convergence is uniform in $n$, we can find a $g_{0}$ such that when $g>g_{0}$, we have $p_{n}(g)<\frac{\delta}{N}$ for $|n|>n_{0}$ and $\sum_{|n|\leq n_{0}}p_{n}(g)<1-\delta$. Similar to the proof of the lemma, the entropy is decreased if we move the $N$ largest $p_{n}(g)$’s in the range $|n|>n_{0}$ to $\frac{\delta}{N}$ while keeping the sum $\sum_{|n|>n_{0}}p_{n}(g)>\delta$ fixed. In this case the entropy is larger than $\delta\log({\frac{N}{\delta}})$, which tends to infinity as $N\rightarrow\infty$.
Obviously the conditions cannot be further relaxed. If $\sum_{n}\tilde{p}_{n}=1$, it is easy to construct examples such that the entropy is bounded as $g\rightarrow\infty$. In an infinite dimensional Hilbert space, entropy in general is not necessarily continuous and there are only a few known restricted continuity properties, e.g. the lower semicontinuity [18]. In mathematical analysis, Tannery’s theorem validates the interchange of infinite sum and limit if the sum is bounded by a fixed convergent sum. In physical situations, such as the thermal partition function at a finite temperature [7], or the entropy of BMN strings at a finite coupling [30], the conditions for Tannery’s theorem are easily verified to be satisfied, so the entropy is a continuous function. However the conditions may not be satisfied asymptotically, as it happens in the infinite temperature limit. So in the case of $\sum_{n}\tilde{p}_{n}=1$, regardless of whether $\tilde{S}:=-\sum_{n}\tilde{p}_{n}\log\tilde{p}_{n}$ is finite, in general it seems undetermined whether $\lim_{g\rightarrow\infty}S(g)$ exists or is equal to $\tilde{S}$.